Consider the following matrix equation.

( "\begin{matrix}
 2 & -1 & 0 \\
 1 & 2 & -1 \\
 0& 1&3
\end{matrix}" ) ( "\begin{matrix}
 x1 \\
 x2 \\
x3
\end{matrix}" ) = ( "\begin{matrix}
 2 \\
 1 \\
3
\end{matrix}" )

Solve the equation by using Thomas algorithm.


1
Expert's answer
2021-11-15T16:11:22-0500

Solution. From the first equation of the system


"2x_1-x_2=2"

Divide through by 2 get


"x_1 -0.5x_2=1."

We write the system of equations as


"\begin{pmatrix}
 1 & -0.5 & 0 \\
 1 & 2 & -1 \\
 0 & 1 & 3 
\end{pmatrix} 
 \begin{pmatrix}
 x_1 \\
 x_2 \\
 x_3 
\end{pmatrix} = \begin{pmatrix}
 1 \\
 1 \\
 3 
\end{pmatrix} ."

From the second equation of the system


"x_1 +2x_2-x_3=1"

Replacing x_1 -0.5*x_2=1 get

"1 +2.5x_2-x_3=1 \to 2.5x_2-x_3=0"

Divide through by 2.5 get


"x_2-\frac{2}{5}x_3=0."

We write the system of equations as


"\begin{pmatrix}
 1 & -0.5 & 0 \\
 0 & 1 & -\frac{2}{5} \\
 0 & 1 & 3 
\end{pmatrix} 
 \begin{pmatrix}
 x_1 \\
 x_2 \\
 x_3 
\end{pmatrix} = \begin{pmatrix}
 1 \\
 0 \\
 3 
\end{pmatrix} ."

From the third equation of the system


"x_2+3x_3=3"

Replacing "x_2-\frac{2}{5}x_3=0" get


"\frac {17}{5} x_3=3"

Divide through by "\frac {17}{5}" get


"x_3= \frac{15}{17}"

We write the system of equations as


"\begin{pmatrix}
 1 & -0.5 & 0 \\
 0 & 1 & -\frac{2}{5} \\
 0 & 0 & 1 
\end{pmatrix} 
 \begin{pmatrix}
 x_1 \\
 x_2 \\
 x_3 
\end{pmatrix} = \begin{pmatrix}
 1 \\
 0 \\
 \frac {15}{17} 
\end{pmatrix} ."

Let us find the roots of the equations of the system

"x_3= \frac{15}{17}"

"x_2=0+\frac{2}{5}x_3 = \frac{2}{5}\times \frac{15}{17}= \frac{6}{17}"

"x_1 = \frac{1}{2}x_2+1=\frac{1}{2} \times \frac{6}{17} +1=\frac{3}{17} + 1 = \frac{20}{17}"

Answer. "x_1 = \frac{20}{17}"; "x_2= \frac{6}{17}"; "x_3= \frac{15}{17}".